Steering Manufacturing produces snow shovels. The selling price per snow shovel is $29.00. There is no beginning inventory. Costs involved in production are: Direct material $5.00 Direct labor 4.00 Variable manufacturing overhead 4.00 Total variable manufacturing costs per unit $13.00 Fixed manufacturing overhead per year $181,170 In addition, the company has fixed selling and administrative costs of $172,400 per year. During the year, Steering produces 54,900 snow shovels and sells 49,560 snow shovels.

What is net income using full costing?

Net income $_______
Answer #1
Direct material 5.00
Direct labor 4.00
Variable manufacturing overhead 4.00
Fixed manufacturing overhead 3.30 =181170/54900
Unit product cost 16.30
Sales revenue 1437240 =49560*29
Less: Cost of goods sold 807828 =49560*16.30
Gross profit 629412
Less: Selling and administrative costs 172400
Net income 457012